OpenStudy (cenaida):

Evaluate 2/5 + (y 1/5) if y = 1/3. A.1/3 B.8/15 C.2/15 D.1/5

OpenStudy (carson889):

Is it supposed to be y multipled by 1/5?

OpenStudy (cenaida):

Im sorry i wrote it wrong its supposed to be like this: Evaluate 2/5 + (y - 1/5) if y = 1/3.

OpenStudy (carson889):

Oh, Well then. For y - 1/5 we must find the common denominator which is 15. I know this because if we multiply 3 and 5. 3: 3,6,9,12,15 and 5: 5,10,15,20. We see that 15 is common. Thus for y = 1/3 we must multiply the numerator and denominator by 5: 1/3 = 5/15. For 1/5 multiply the numerator and denominator by 3: 1/5 = 3/15. Now we know: 2/5+ (5/15 - 3/15) = 2/5 + 2/15. Now 2/5 multiplied by 3 in both the numerator and denominator is 6/15. So we can finish the question: 6/15 + 2/15 = 8/15. Voila! The answer is B

OpenStudy (cenaida):

Thanks! great answer :)

OpenStudy (anonymous):

8/15